OpenStudy (anonymous):

How do i solve for x and express my answer in simplest answer form? (4/x) - (3/(x+1)) = 7

OpenStudy (anonymous):

please vote http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

i will help u

OpenStudy (anonymous):

(4/x) - (3/x + 1) = 7 1/x - 1 = 7 1/x = 8 x = 1/8 Answer: 1/8

OpenStudy (anonymous):

that is the answer and i solved it please vote

OpenStudy (anonymous):

did u vote

OpenStudy (anonymous):

@Nick_B

OpenStudy (anonymous):

i got x= 1/8 however, the choices are different. I have (-3+/- Square root 37) /7

OpenStudy (anonymous):

@katragaddasaichandra

OpenStudy (anonymous):

did u vote

OpenStudy (darkigloo):

First you have to find the common denominator. Did you do that?

OpenStudy (anonymous):

Yeah x(x+1)

OpenStudy (darkigloo):

right. \[\frac{ 4(x+1) }{ (x)(x+1) } - \frac{ 3x }{ )x)(x+1)} = \frac{ 7(x)(x+1) }{ (x)(x+1) }\] Now since they are all under the same denominator, you can solve it like this: \[4(x+1) - 3x = 7(x)(x+1)\]

OpenStudy (radar):

Another hint, it ain't 1/8

OpenStudy (radar):

\[4x + 4 -3x = 7x ^{2} +7x\]

OpenStudy (radar):

Maybe it won't factor nicely, Use the quadratic formula to solve for the irrational roots.